OpenStudy (anonymous):

determine whether y varies directly with x. If so, find the constant of variation k and write the equation X 6 18 54 162 Y 24 72 216 648 Options A-yes, K=4;y=4x B-yes,K=3;y=3x C-yes,K=6;y=6x D-no Please show work Please help me

OpenStudy (unklerhaukus):

what is the ratio of the corresponding terms in each coordinate ,

OpenStudy (unklerhaukus):

X Y 6 24 18 72 54 216 162 648 \[(x_1,y_1)=(6,24)\]\[(x_2,y_2)=(18,72)\]\[\dots\] \[(x_n,y_n)=(x_n,\alpha x_n)\] \[y(x)\propto x\] \[y(x)=\alpha x\]

OpenStudy (anonymous):

i need help please help me

OpenStudy (anonymous):

What exactly is the type of mathematics are you doing? Algebra 2? Pre-Calc?

OpenStudy (anonymous):

algebra 2

OpenStudy (anonymous):

XD.... sorry..... I'm only 13, and I am completing geometry this summer so I can skip ahead to Algebra 2 in about 3 1/2 weeks. Sorry, but I can't help you in this situation.... I apologize for this.....

OpenStudy (anonymous):

maybe someone else of a higher education could explain, or at least, brief us on this problematic conundrum....

OpenStudy (unklerhaukus):

a point is \((x_1,y_1)=(6,24)\) \[y_1(x_1)=\alpha x_1\] what is alpha/

OpenStudy (anonymous):

i dont kno

OpenStudy (unklerhaukus):

\[24=\alpha\times6\]

OpenStudy (anonymous):

thanks for the medal anikate

OpenStudy (anikate):

no problem, it's a good question

OpenStudy (anikate):

@sami-21

OpenStudy (anikate):

@Agent_Sniffles

OpenStudy (anikate):

sry gtg my mom wants me to vacum the living room

OpenStudy (anikate):

@AccessDenied

OpenStudy (anonymous):

its ok

OpenStudy (anikate):

very sry I couldnt help you

OpenStudy (anonymous):

np

OpenStudy (anikate):

try to look up youtube vides they really helped me

OpenStudy (anonymous):

i tried that already

OpenStudy (anikate):

ohh... then sry I dont know what to say :(

OpenStudy (anikate):

try posting the question in an easier understandable way?

OpenStudy (anonymous):

but that is the only way i can

OpenStudy (anonymous):

@hereweg0

OpenStudy (anonymous):

@Bambi @SmoothMath @timmtamm96 i need help

OpenStudy (unklerhaukus):

please read what i have posted above @tabbycat737

OpenStudy (anonymous):

i kno i read that but i dont understand the rest

OpenStudy (unklerhaukus):

iff y varies directly with x we can wright \[y\propto x\]\[y=\alpha x\] where alpha is the constant of variation

OpenStudy (unklerhaukus):

if y varies directly with x \[y=αx\] \[y_1=\alpha x_1\]\[y_2=\alpha x_2\]etc

OpenStudy (unklerhaukus):

the table shows \(x_n\)'s and their corresponding \(y_n\)'s, try subbing in a point \((x,y)\) for the table to the direct variation equation

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

still lost with this one

OpenStudy (unklerhaukus):

what are some points in the function?

OpenStudy (anonymous):

ther are 18 and 72 ther is 54 an 216 ther is 162 and 648